# Design Principles of Zero-Shot Self-Supervised Unknown Emitter Detectors

### v1

@@ Abstract
The proliferation of wireless devices necessitates more robust and reliable emitter detection and identification for critical tasks such as spectrum management and network security. Existing studies exploring methods for unknown emitters identification, however, are typically hindered by their dependence on labeled or proprietary datasets, unrealistic assumptions (e.g. all samples with identical transmitted messages), or deficiency of systematic evaluations across different architectures and design dimensions. In this work, we present a comprehensive evaluation of unknown emitter detection systems across key aspects of the design space, focusing on data modality, learning approaches, and feature learning modules. We demonstrate that prior self-supervised, zero-shot emitter detection approaches commonly use datasets with identical transmitted messages. To address this limitation, we propose a 2D-Constellation data modality for scenarios with varying messages, achieving up to a 40% performance improvement in ROC-AUC, NMI, and F1 metrics compared to conventional raw I/Q data. Furthermore, we introduce interpretable Kolmogorov–Arnold Networks (KANs) to enhance model transparency, and a Singular Value Decomposition (SVD)-based initialization procedure for feature learning modules operating on sparse 2D-Constellation data, which improves the performance of Deep Clustering approaches by up to 40% across the same metrics comparing to the modules without SVD initialization. We evaluate all data modalities and learning modules across three learning approaches: Deep Clustering, Auto Encoder and Contrastive Learning.
@@ I Introduction
The management and operation of wireless communication networks increasingly depend on situational and spectrum awareness to dynamically adapt to environmental variations and user-imposed demands. Spectrum awareness and network security are envisioned as part of the main research directions [polese2023understanding] for the future Open Radio Access Networks (O-RAN) [alliance2020ran], but the same techniques can also be used in interference management, regulatory compliance and spectrum optimization. Together with the development of modulation and technology classification [milosheski2023self], they contribute to the creation of efficient, spectrum-aware, wireless communication networks [ouyang2023integrated]. One of the key challenges related to spectrum awareness lies in the reliable detection and identification of emitters within the communication range, on the conceptual level depicted in Figure 1(b), which is relying on robust signal processing techniques and increasingly leveraging advanced machine learning approaches. In this respect, the research community distinguishes between two complementary processes within the broader context of radio signal intelligence and spectrum awareness, the Specific Emitter Identification (SEI) and the Unknown Emitter Detection (UED).
SEI focuses on recognizing and classifying known emitters under a closed-set assumption based on unique, device-specific signal alterations—such as transient behaviors, hardware-induced imperfections, or modulation patterns—allowing precise attribution of transmissions to individual sources [ding2018specific]. This makes SEI particularly well suited for application in network security [robinson2020dilated] with the aim to distinguish emitters that are allowed to operate in a certain wireless network, in which case they are transmitting useful information and can be referred to as transmitters. In contrast, UED is working under open-set assumption and it aims to identify signals that deviate from any known emitter profile, detecting and isolating novel, spoofed, or adversarial sources that fall outside the established feature space of the known emitters. In network security application, UED is used to detect and reject or flag signals to prevent their misclassification by SEI. As such, these processes form a closed-loop detection and identification system whereby SEI provides fine-grained, instance-level identification within a known feature space used for training, whereas UED ensures robustness and adaptability by flagging outliers and prompting model updates or human review for inclusion of novel devices, together maintaining situational awareness and security in dynamic or contested radio environments.
@@
(a) Design space for building deep learning model for detection and identification of emitters.
(b) The concept of detection and identification of emitters.
@@
Figure 1: The concept and design space for detection and identification of known/unknown emitters.
@@
The approach to developing a model capable of learning to distinguish emitters based on device-specific signal alterations largely depends on the application and systematic consideration of all the existing design choices, as visualized in Figure 1(a). Most of the existing approaches are focusing on closed-set classification problem. Such approaches consider that all the emitters are known and labeled transmission data is available for each of them for developing supervised classification models [wong2018clustering, han2022proceed], reflecting the authentication use case of known devices. Some of these models achieved remarkable performance in distinguishing known devices [luo2022transformer], [basak2021combined], [wang2021specific], even when thousands of such devices [robinson2020dilated] exist in the network. However, in real-world operating scenarios with the rising volume and types of wireless devices, new transmitters could appear regularly in the network, which could pose significant security and management challenges. In such open-set setup there is little or no labeled data available for training, making the supervised learning for model development unsuitable. The open-set challenges are divided into two distinct subsets: few-shot learning [xu2023few, huang2022deep, liu2023overcoming] and zero-shot learning [robinson2021novel]. Few-shot learning follows the supervised evaluation protocol but operates under severe label constraints, with the number of labeled samples ranging from just a few (1–20) [huang2022deep] to several hundred [peng2023supervised]. In contrast, zero-shot learning aims to detect emitters that are entirely absent during training [dong2021sr2cnn, wen2022drsn].
Self-supervised learning (SSL) is a machine learning paradigm capable of learning latent feature representations from large volumes of unlabeled data. By formulating and solving pretext tasks derived directly from the data, SSL enables the development of models that generalize well without relying on extensive manual labeling. This makes it particularly well-suited for open-set detection and identification of emitters [peng2023supervised, stankowicz2021unsupervised, zhang2023transmitter], closely aligning with realistic deployment conditions, where the diversity of emitters and the dynamic nature of the spectrum make comprehensive labeling impractical.
@@
Learning module is the core part of the learning process, which depends on the application constrains, such as available hardware, inference time requirements, etc. Thus, it directly influences many of the other design choices, such as performance, complexity, explainability, and interpretability. The existing model architectures, already proven in related domains, which include Convolutional Neural Networks (CNN) [robinson2021novel], transformers [luo2022transformer], Long Short-Term Memory (LSTM) recurrent neural networks [al2021deeplora] and/or combinations with Multilayer Perceptron (MLP) [stankowicz2021unsupervised], usually provide decent performance [robinson2021riftnet]. However, designing a dedicated model for the radio signal specifics and predefined use case can lead to significant complexity reduction [milosheski2024spectrum]. The choice of a model architecture also affects its explainability and interpretability. For the deep learning models, there are several explainability approaches, such as Local Interpretable Model-agnostic Explanations (LIME) [ribeiro2016should], SHapley Additive exPlanations) (SHAP) [lundberg2017unified], and saliency maps [simonyan2013deep], whereas the issue of model interpretability remains largely unaddressed. This is primarily due to the complexity and opacity of the deep learning modules employed, typically large CNNs or self-attention-based architectures [luo2022transformer], which make it difficult to understand or explain model decisions, while also lacking interpretability. So far, the interpretable models are the classic machine learning (ML) approaches, such as Support Vector Machine (SVM), and logistic regression [riyaz2018deep], which are significantly outperformed by the deep learning models, while interpretability is overlooked in most of the recent works. However, due to the potential use in security-related applications, such as device authentication, interpretability of the detection and identification of emitters remains a crucial design choice. To address this limitation, we include in our study the Kolmogorov–Arnold Network (KAN) [moradi2024kolmogorov] as a feature learning module in SSL framework, an inherently interpretable neural network architecture that achieves comparable performance to the standard deep learning models.
In this work, we conduct a comprehensive evaluation across multiple zero-shot SSL methods, carefully considering their specific design choices, and two WiFi datasets [sankhe2019oracle, hanna_wisig_2022], which allow investigating the impact of transmitted message structure and content assuming same and different messages on the robustness of detection and identification of emitters. We summarize our contributions as follows:
@@
Ref. / Work Problem Setup Data Modality ML Approach Feature Extractor Interpretability Level Supervision Setup Results / Notes
[hao2023contrastive] Self-Supervised SEI Raw I/Q CL CNN DMM level Requires the number of training emitters 93% F1
[zhang2023transmitter] Supervised Zero-Shot UED Raw I/Q CL CNN DMM level Embedded in contrastive loss 80% ACC
@@
[zhan2023mcrff] Meta-CL for SEI Raw I/Q Meta-CL CNN DMM level Embedded in contrastive loss 83% ACC
@@
[zha2023cross] Few-Shot SEI Raw I/Q SimSiam Contrastive Learning CNN None Few-shot fine-tuning 80% ACC
@@
Proposed Self-Supervised Zero-Shot UED Raw I/Q, 2D-Constellation CL, DC, AE CNN, KAN DMM and FE levels Unsupervised 85% F1, 94% ACC
@@
  • We define a generic workflow for zero-shot self-supervised detection and identification of emitters, aligned with the design space depicted in Fig. 1(a), and propose a new method that is aligned with this workflow and relies on an interpretable feature extractor from raw IQ samples.
  • We propose a 2D-Constellation data modality for zero-shot self-supervised UED which shows up to 40% performance improvement compared to conventional raw I/Q input format under different-messages scenario. To the best of our knowledge, we are the first to address the challenge of different messages in Self-Supervised SEI/UED.
  • We show that the KAN-based interpretable feature learning module for zero-shot self-supervised UED demonstrates same performance or better than conventional CNN architecture.
  • We propose a singular value decomposition (SVD) initialization procedure for feature learning modules that improves the performance up to 40% compared to random initialization using sparse 2D-Constellation input format.
  • We demonstrate that Deep Clustering approach produces the best unknown emitter detection performance under different-messages scenario.
This paper is organized as follows. Section II summarizes the related work. Section III provides the background and problem statement, while Section IV elaborates on the design choices of the emitter detection and identification framework. Section V introduces the methodology and Section VI provides a thorough analysis of the results. Section VII concludes the paper.
@@ II Related work
When focusing on semi-supervised and unsupervised deep learning approaches for detection and identification of emitters the number of related studies is rather low. Table I summarizes the related works focused on SEI and UED tasks based on semi-supervised and unsupervised deep learning approaches, structured along the design space presented in Fig. 1(a). As shown, prior works are predominantly focusing on few-shot self-supervised SEI using raw I/Q data with Contrastive Learning (CL) [wu2023specific, zha2023cross, liu2023overcoming] or Auto Encoder (AE) [huang2022deep, xu2023few, yao2023few], with CNN-based feature extractors being the most common backbone. Besides the necessity for labels, the interpretability of these approaches, if supported at all, is limited to the decision-making stage, often via distance-based metrics, while their feature extractors remain opaque. There is also only one work for Self-Supervised SEI utilizing 2D Constellation Data Modality [liu2022specific], which is also designed for Few-Shot set-up limiting it’s applications. Moreover, they don’t either theoretically or experimentally demonstrate the need of this transformation. In this work, we investigate the Desing Space of approaches for Self-Supervised UAD ny usage of labels during the training - a more challenging problem. We prove theoretically and experimentally that in complete labless case the scenarios of signals transmission greatly affects the right design choice. We show the necessity of using 2D Constellation Data Modality when transmitted messages are different.
@@
Despite differences in methodology, only two studies [krasnov2025novel, hao2023contrastive] are fully unsupervised, requiring no labels at any stage. In [hao2023contrastive], authors demonstrate an F1 score of $93\%$ and over 98% accuracy in a closed-set SEI task using a contrastive learning framework, underscoring the effectiveness of the SSL approaches in overcoming the labeling bottleneck. However, this work uses a proprietary non-public dataset, which makes the results not replicable. Authors of [krasnov2025novel], on the other hand, use publicly available WiSig dataset [hanna_wisig_2022], but it only contains signals transmitting the same messages, making the setup less realistic. They achieve an F1 score between $80\%$ and $87\%$ on self-supervised UED task depending on the number of emitters in the training set versus all emitters in the testing set. In addition, mentioned approaches still lack intepretability of the features learning module.
@@
In this work, we aim to fill the identified gaps by conducting experiments on two publicly available datasets [hanna_wisig_2022, sankhe2019oracle] supporting investigation of the impact of transmitting same and different messages. We are focusing on a zero-shot self-supervised open-set UED task and in contrast to related works we introduce interpretability both at the embedding and feature-extractor levels, providing the first systematic performance-versus-interpretability comparison across Deep Clustering (DC), Contrastive Learning (CL), and Auto Encoder (AE) paradigms.
@@ III-A RF Fingerprint
In this work, we focus on detection and identification of emitters using Orthogonal Frequency Division Multiplexing (OFDM) modulation scheme, which is widely spread and commonly used in wireless communications including in WiFi. In this modulation scheme, each transmitted symbol corresponds to a complex-valued point in the I/Q (In-phase/Quadrature) constellation diagram, which represents the signal’s amplitude and phase characteristics. Figure 2 shows an example constellation diagram of the OFDM modulation scheme with the axes representing the I and Q components of signal points. Each OFDM signal consists of clean modulated message and distortions caused by hardware imperfections and channel noise. While channel noise is affected by environment conditions, hardware imperfections, often referred as radio frequency (RF) fingerprint, are caused by tiny imperfections of individual emitter analog components, making them emitter specific. In this study, we explore these hardware imperfections for distinguishing known and unknown emitters. Simplifying, RF Fingerprint [xie2024radio] can be modeled as:
@@
where $\Delta f$ and $\phi(t)$ are frequency and phase offsets in signal $x(t)$ caused by hardware imperfections. Treating them as small perturbations, we apply first order Taylor decomposition to divide signal into clean and corrupted components:
@@
where $x_{symbols}(t)$ represents clean signal without hardware imperfections and $r_{emitter}(t)$ stands for the RF Fingerprint part. To illustrate the distinctions created by the fingerprint $r_{emitter}(t)$ for two measured transmissions of the same symbols, we plot Figure 2. The figure demonstrates the two signals from two different emitters from the WiSig Dataset [hanna_wisig_2022] where the yellow dots represent Emitter 1 and black dots represent Emitter 2.
The Physical (PHY) layer of an 802.11 frame that is responsible for transmitting data over the wireless medium by modulating it into a signal consists of a preamble, a header and a payload. The preamble and header have fixed length, whereby the first has fixed repeating sequence of bits used for synchronization and the second contains varying control information like the transmission rate and payload length. These are followed by a varying payload of the length specified in a header containing a medium access control (MAC) frame with the actual data and a trailer with error-checking information. In this work, we define two signal receiving scenarios: Same Message Scenario for the cases when the preamble is sensed and Random Message Scenario.
III-B Same Messages Scenario (SM)
If only looking at a preamble, a signal with the same content is transmitted by every emitter in the network. The preamble signal coming from $i^{\textbf{th}}$ emitter, using Eq. 2, can be described as:

### v2

@@ Abstract
Robust situational awareness in contested wireless environments requires the ability to detect unauthorized or hostile emitters without prior knowledge of their signatures. Existing studies on Unknown Emitter Detection (UED) and identification are hindered by their reliance on labeled or proprietary datasets, unrealistic assumptions (e.g., all samples containing identical transmitted messages), or a lack of systematic evaluation across architectures and design dimensions. In this work, we identify the design dimensions of machine learning based UED and introduce a structured workflow for developing such systems. The workflow, suitable for guiding studies and end-to-end label-free UED designs, consists of four key components: (a) data modality, (b) feature learning module, (c) machine learning approach, and (d) decision-making module. Using this workflow, we examine the design principles of UED in two distinct transmission scenarios: same messages scenario (SMS) and different messages scenario (DMS). Our investigation yields several key findings. (a) Under realistic DMS, the 2D constellation data modality achieves up to a 20-percentage-point improvement in ROC-AUC compared to the conventional raw I/Q representation. (b) Kolmogorov-Arnold Networks (KANs) provide interpretable representations while achieving performance comparable to CNNs in both scenarios (c) for DMS, the best-performing configuration combines SVD-initialized KANs with a deep clustering approach and the 2D constellation modality, improving ROC-AUC by up to 20 percentage points over standard KANs with identical workflow components. (d) through analysis of the decision-making module, we determine the optimal number of clusters for environments with varying numbers of known emitters.
@@ I Introduction
The management and operation of wireless communication networks increasingly depend on situational and spectrum awareness to dynamically adapt to environmental variations and user/application-imposed demands. Spectrum awareness and network security are envisioned as part of the main research directions [polese2023understanding] for the future Open Radio Access Networks (O-RAN) [alliance2020ran], but the same techniques can also be used in interference management, regulatory compliance and spectrum optimization. Together with the development of modulation and technology classification [milosheski2023self], they contribute to the creation of efficient, spectrum-aware, wireless communication networks [ouyang2023integrated]. One of the key challenges related to spectrum awareness lies in the reliable detection and identification of emitters within the communication range, on the conceptual level depicted in Figure 1(a). It relies on robust signal processing techniques and increasingly leverages advanced machine learning approaches. In this respect, the research community distinguishes between two complementary processes within the broader context of devices recognition: the Specific Emitter Identification (SEI) and the Unknown Emitter Detection (UED).
(a) The concept of self-supervised detection and identification of emitters.
@@
(b) Design space for building deep learning model for detection and identification of emitters.
@@
Figure 1: The design space for ML detection and identification of known/unknown emitters.
SEI focuses on recognizing and classifying known emitters under a closed-set assumption based on unique, device-specific signal alterations such as transient behaviors, hardware-induced imperfections, or modulation patterns allowing precise attribution of transmissions to individual sources [ding2018specific]. This makes SEI particularly well suited for network security applications [robinson2020dilated]. The aim is to distinguish emitters that are allowed to operate in a certain wireless network, in which case they are transmitting useful information and can be referred to as transmitters. In contrast, UED [apfeld2021recognition] is working under open-set assumption and it aims to identify signals that deviate from any known emitter profile, detecting and isolating novel, spoofed, or adversarial sources that fall outside the established feature space of the known emitters. In network security application, UED is used to detect and reject or flag unrecognized signals to prevent their misclassification by SEI. As such, these processes form a closed-loop detection and identification system whereby SEI provides fine-grained, instance-level identification within a known feature space used for training, whereas UED ensures robustness and adaptability by flagging outliers and prompting model updates or human review for inclusion of novel devices, together maintaining situational awareness and security in dynamic or contested radio environments.
@@
Most of the existing approaches are focusing on the closed-set classification problem. Such approaches consider that all the emitters are known and labeled transmission data is available for each of them for developing supervised classification models [wong2018clustering, han2022proceed], reflecting the authentication use case of known devices. Some of these models achieved remarkable performance in distinguishing known devices [luo2022transformer, basak2021combined, wang2021specific], even when thousands of such devices [robinson2020dilated] exist in the network. However, in real-world operating scenarios with the rising volume and types of wireless devices, new transmitters could appear regularly in the network, which could pose significant security and management challenges. In such open-set setup there is little or no labeled data available for training, making the supervised learning for model development unsuitable. The open-set challenges are divided into two distinct subsets: few-shot learning [xu2023few, huang2022deep, liu2023overcoming] and zero-shot learning [robinson2021novel]. Few-shot learning follows the supervised evaluation protocol but operates under severe label constraints, with the number of labeled samples ranging from just a few (1–20) [huang2022deep] to several hundred [peng2023supervised]. In contrast, zero-shot learning aims to detect emitters that are entirely absent during model training [dong2021sr2cnn, wen2022drsn].
We notice that, overall, the approach to developing a system capable of learning to distinguish emitters based on device-specific signal alterations largely depends on the application and systematic consideration of all the existing design choices, as visualized in Figure 1(b). The problem set-up considering zero-shot and few shot learning is one dimension. Other dimensions relate to data modality, feature learning, ML approach, etc. The selection of data modality may significantly improve the final performance, but it also influences the complexity of the feature learning module. The design of the feature learning module influences many of the other design choices, such as performance, complexity, explainability, and interpretability. The existing model architectures such as Convolutional Neural Networks (CNN) [robinson2021novel], transformers [luo2022transformer], Long Short-Term Memory (LSTM) recurrent neural networks [al2021deeplora] and/or combinations with Multilayer Perceptron (MLP) [stankowicz2021unsupervised] have been already proven in related domains and they usually provide satisfactory performance [robinson2021riftnet]. However, designing a dedicated model for the radio signal specifics and predefined use cases can lead to significant complexity reduction [milosheski2024spectrum].
@@
The choice of the neural architecture for feature learning also affects its explainability and interpretability. For the deep learning models, there are several explainability approaches, such as Local Interpretable Model-agnostic Explanations (LIME) [ribeiro2016should], SHapley Additive exPlanations) (SHAP) [lundberg2017unified], and saliency maps [simonyan2013deep], whereas the issue of model interpretability remains largely unaddressed. This is primarily due to the complexity and opacity of the deep learning modules employed, typically including large CNNs or self-attention-based architectures [luo2022transformer], which make it difficult to understand or explain model decisions, while also lacking interpretability. So far, the interpretable models are the classic machine learning (ML) approaches, such as Support Vector Machine (SVM), and logistic regression [riyaz2018deep], which are significantly outperformed by the deep learning models, while interpretability is overlooked in most of the recent works. However, due to the potential use in security-related applications, such as device authentication, interpretability of the detection and identification of emitters remains an important design choice.
For zero-shot set-ups, Self-Supervised Learning (SSL) [liu2021self] techniques are the most suitable ML approaches. SSL represents a family of machine learning methods, which create supervision signal without using ground truth labels but through data manipulation techniques such pseudo-labeling, augmentations, reconstruction etc. By formulating and solving pretext tasks derived directly from the data, SSL enables the development of models that generalize well without relying on manual labeling. This makes it particularly well-suited for open-set detection and identification of emitters [peng2023supervised, stankowicz2021unsupervised, zhang2023transmitter], closely aligning with realistic deployment conditions, where the diversity of emitters and the dynamic nature of the spectrum make comprehensive labeling impractical.
@@
Ref. Problem Setup Data Modality ML Approach Feature Extractor Interpret / Explain Supervision Setup Results / Notes
[hao2023contrastive] Self-supervised SEI/UED Raw I/Q CL CNN DMM level No. of training emitters 93% F1
@@
[zha2023cross] Few-Shot SEI Raw I/Q SimSiam CL CNN None Few-shot fine-tuning 80% ACC
[sun2025few] Few-Shot SEI Raw I/Q CL CNN None Few-shot fine-tuning 95% ACC
@@
Ours Self-supervised zero-shot UED Raw I/Q, 2D constellation CL, DC, AE CNN, KAN DMM and FE levels Unsupervised 85% F1, 94% ACC
@@
In this work, we identify the design dimensions of machine learning based UED as depicted in Figure 1(b): 1) problem set-up, 2) data modality, 3) feature learning, 4) ML approach, 5) decision making, 6) evaluation metrics and 7) explainability and interpretability. Focusing on realistic environments where no prior information is available, we consider a zero-shot set-up that inherently requires a SSL approach. Consequently, we introduce a structured workflow for developing such systems. The workflow, suitable for guiding studies for end-to-end label-free UED designs. Using this workflow, we examine the design principles of UED in two distinct transmission scenarios: same messages scenario (SMS) and different messages scenario (DMS) enabled by two distinct datasets [sankhe2019oracle, hanna_wisig_2022]. Throughout the study we consider two possible data modalities, three different feature learning and three different SSL approaches evaluated through several metrics including also explainability and interpretability considerations. For replicability purposes, the scripts are available as open source [^1] https://github.com/sensorlab/ZeroUED. We summarize our contributions as follows:
  • We formalize a structured, self-supervised workflow for zero-shot UED on unlabeled data, consisting of four key components: a) data modality, b) feature learning module, c) ML approach, and d) decision-making module.
  • Through the lens of the proposed workflow, we investigate the design principles of designing a an UED on two transmission scenarios: Same Messages Scenario (SMS) and Different Messages Scenario (DMS), two data modalities, two neural architectures, including KAN designed for interpretability, and three SSL ML approaches.
  • During our investigation, we find the following. a) Under the realistic DMS scenario, 2D constellation data modality shows up to 20 p.p. ROC-AUC metric improvement compared to conventional raw I/Q representation. b) Kolmogorov-Arnold networks (KAN s) are interpretable and have performance comparable to CNNs, a convenient alternative that suffers from ”black-box” design, across SMS and DMS. c) We observe that for DMS, the best performing configuration is SVD initialized KAN with deep clustering approach and 2D constellation data modality. This configuration improves ROC-AUC performance up to 20 p.p. compared to the standard KAN with the same other parts of the workflow in DMS. d) Investigating the decision-making module, we determine the optimal number of clusters for environments with different numbers of known emitters.
This rest of the paper is organized as follows. Sec. II summarizes the related work, Sec. III provides the background and problem statement, while Sec. IV elaborates on the design choices of the emitter detection and identification framework. Sec. V describes the evaluation methodology and Sec. VI analyzed the results. Finally, Sec. VII concludes the paper.
@@ II Related work
When focusing on semi-supervised and unsupervised deep learning approaches for detection and identification of emitters the number of related studies is rather low. Table I summarizes the related works focused on SEI and UED tasks based on semi-supervised and unsupervised deep learning approaches, structured along the design space presented in Figure  1(b). As shown, prior works are predominantly focusing on few-shot self-supervised SEI using raw I/Q data with Contrastive Learning (CL) [wu2023specific, zha2023cross, liu2023overcoming, sun2025few] or Auto Encoder (AE) [huang2022deep, xu2023few, yao2023few], with CNN-based feature extractors being the most common backbone. Besides the necessity for labels, the interpretability of these approaches, if supported at all, is limited to the decision-making stage, often via distance-based metrics, while their feature extractors remain opaque. There is also only one work for self-supervised SEI utilizing 2D constellation data modality [liu2022specific], which is also designed for few-shot set-up limiting its applications, whereby the authors do not demonstrate either theoretically nor experimentally the need of this transformation. In this work, we investigate the design space of approaches for self-supervised UED without using labels during the training, which is a more challenging problem. We prove theoretically and experimentally that in completely label-less case the scenarios of signal transmission greatly affect the correct design choice. We show the advantage of using 2D constellation data modality when transmitted messages are different.
@@
Despite differences in methodology, only two studies [krasnov2025novel, hao2023contrastive] are fully unsupervised, requiring no labels at any stage. In [hao2023contrastive], as part of the study, authors demonstrate 91% accuracy in a zero-shot UED task using a contrastive learning framework, underscoring the effectiveness of SSL approaches in overcoming the labeling bottleneck. However, this work uses a proprietary non-public dataset, which makes the results non-replicable. Authors of [krasnov2025novel], on the other hand, use publicly available WiSig dataset [hanna_wisig_2022], but it only contains signals transmitting the same messages, making the setup less realistic. They achieve an F1 score between $80\%$ and $87\%$ on self-supervised UED task depending on the number of emitters in the training set versus all emitters in the testing set. All above-mentioned approaches lack interpretability of the features learning module.
@@
In this work, we aim to fill the identified gaps by conducting experiments on two publicly available datasets [hanna_wisig_2022, sankhe2019oracle] supporting also the investigation of the impact of transmitting same and different messages. We are focusing on a zero-shot self-supervised open-set UED task and in contrast to related works we introduce interpretability both at the embedding and feature-extractor levels, providing the first systematic performance-versus-interpretability comparison across Deep Clustering (DC), Contrastive Learning (CL), and Auto Encoder (AE) paradigms.
Figure 2: Yellow dots represent an arbitrary signal from the first emitter from the WiSig Dataset [hanna_wisig_2022], and black dots stand for the arbitrary signal from the second Emitter. The signals carry the same message and differ only by environmental noise and hardware imperfections.
Figure 3: Formal self-supervised ML based UED workflow.
@@ III-A RF Fingerprint
In this work, we focus on detection and identification of emitters using Orthogonal Frequency Division Multiplexing (OFDM) modulation scheme. In this modulation scheme, each transmitted symbol corresponds to a complex-valued point in the I/Q (In-phase/Quadrature) constellation diagram, which represents the signal’s amplitude and phase characteristics. Figure 2 shows an example constellation diagram of the OFDM modulation scheme with the axes representing the I and Q components of signal points. Each OFDM signal consists of clean modulated message and distortions caused by hardware imperfections and channel noise. While channel noise is affected by environment conditions, hardware imperfections, often referred as radio frequency (RF) fingerprint, are caused by tiny imperfections of individual emitter analog components, making them emitter specific. In this study, we explore these hardware imperfections for distinguishing known and unknown emitters. Simplifying, RF fingerprint [xie2024radio] can be modeled as:
@@
where $\Delta f$ and $\phi(t)$ are frequency and phase offsets in signal $x(t)$ caused by hardware imperfections. Treating them as small perturbations, we can apply first order Taylor decomposition to divide signal into clean and corrupted components:
@@
where $x_{symbols}(t)$ represents clean signal without hardware imperfections and $r_{emitter}(t)$ stands for the RF fingerprint part. Figure 2 illustrates the distinctions created by the fingerprint $r_{emitter}(t)$ for two measured transmissions of the same symbols from the WiSig Dataset [hanna_wisig_2022] by two different emitters, where the yellow dots represent Emitter 1 and the black dots represent Emitter 2.
Clearly, the challenge in distinguishing $r_{emitter}(t)$ in the case of known transmitted messages differs from the case of more general unknown transmitted messages, as we show in this work by considering two datasets. The physical (PHY) layer of transceiver is responsible for sending the data over the wireless medium by modulating it onto a carrier signal. It consists of a preamble, a header and a payload. The preamble and header have fixed length, whereby the first also has a fixed repeating sequence of bits used for synchronization while the second contains varying control information like the transmission rate and payload length. These are followed by a varying payload of the length specified in a header containing a medium access control (MAC) frame with the actual data and a trailer with error-checking information. Taking this into account, we consider in this work two signal receiving scenarios: Same Messages Scenario for the cases when only the preamble symbols are considered and Different Messages Scenario when all received symbols are considered.
III-B Same Messages Scenario (SMS)
If only looking at a preamble, a signal with the same content is transmitted by every emitter in the network. The preamble signal coming from the $i^{\textbf{th}}$ emitter, using Eq. 2, can be described as: